If the dimensions of the length are expressed as Gxcyhz; where G, c and h are the universal gravitational constant, speed of light and Planck’s constant respectively, then


 


(a)     X = ½, y = ½


(b)     x = 1/2 , z = 1/2


(c)     y = ½, z = 3/2


(d)     y = –3/2, z = ½

Grade:upto college level

1 Answers

Avinash
8 years ago
L = [M-1L3T-2]x [LT-1]y [ML2T-1]z
should has zero as power of M, so x = z.
Sox = 1/2 , z = 1/2